OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 1) and (–2, 5)? A. -5/4 B. -4/5 C. 5/4 D. 4/5 I got A is that right?

OpenStudy (anonymous):

Not correct. Slope is \[m=\frac{rise}{run}=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

You have your x's and y's switched.

OpenStudy (anonymous):

oh, so if I switched them, would it make it 5/4?

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (3,1) x1 = 3 and y1 = 1 (-2,5) x2 = -2 and y2 = 5 now sub slope(m) = (5 - 1) / (-2-3) slope(m) = 4/-5 slope -4/5 do you understand ?

OpenStudy (anonymous):

Yes, Thank you for explaining it :)

OpenStudy (texaschic101):

sure thing...glad to help :)